OpenStudy (anonymous):

Write a single rule that combines these two translations (x, y) to (x - 7, y) (x, y) to (x + 14, y + 10)

hartnn (hartnn):

new X=x-7 newest X'=X+14 so X'=x-7+14=x+7 got this? can u do the same for y?

OpenStudy (anonymous):

i get confused when you substitute

hartnn (hartnn):

so lets take a value: say old x=4 what would it be at the end of translation? first x-7 so 4-7=-3 then x+14 so -3+14=11 so old was 4,newest=11 so the translation is x+7 got this?

OpenStudy (anonymous):

yes i got that

hartnn (hartnn):

now can u do same thing for y ?

OpenStudy (anonymous):

y+10

hartnn (hartnn):

yup,correct :)

OpenStudy (anonymous):

thankyou for helping

hartnn (hartnn):

so the single rule would be (x+7,y+10)

hartnn (hartnn):

welcome :)